# Landau.bib

### v1

@@ 2.3 Assumptions on curves
For the hard potential cases in Assumption 2.2, we assume $f\in L^{\infty}([0,T];L^{1}_{0,2+\gamma}(\mathbb{R}^{3}\times\mathbb{R}^{3}))$ and $f\in L^{\infty}([0,T];L^{p}_{v}L^{1}_{x})$ for some $\frac{3}{3-\gamma}<p$, which ensures that
$\langle v\rangle^{\gamma}f\in L^{\infty}([0,T];L^{p^{\prime}}_{v}L^{1}_{x})\quad\forall\,\max\{\frac{3}{3-\gamma},\frac{p\gamma}{2+\gamma}\}<p^{\prime}<p.$
Indeed, by Hölder inequality, we have
$\|\langle v\rangle^{|\gamma|}f_{t}\|_{L^{p^{\prime}}_{v}L^{1}_{x}}\leq\|f_{t}\|_{L^{p}_{v}L^{1}_{x}}^{\frac{p-p^{\prime}}{p}}\|f_{t}\|_{L^{1}_{0,2+|\gamma|}(\mathbb{R}^{3}\times\mathbb{R}^{3})}^{\frac{p^{\prime}}{p}}.$
In the follow-up Part II paper, we will show the existence and regularity of the fuzzy Landau equation (1.1).
The variational characterisation Theorem 1.1 for fuzzy Landau equations (1.1) with kernels satisfying Assumption 2.1 are restricted to dimension three, but we expect that the solvability results in Part II are applicable to all dimensions (with appropriate ranges of $\gamma$).
Concerning the propagation of moments and $L^{p}_{v}L^{1}_{x}$ norms, we first see the case of the spatial kernel $\kappa=1$. We integrate (1.1) over $x\in\mathbb{R}^{3}$ and define $F(v)=\int_{\mathbb{R}^{3}}f{\,\rm d}x$, which satisfies the homogeneous Landau equations (2.6). The regularity results for homogeneous Landau equations hold: For $\gamma\in[-2,1]$, all moments and $L^{p}$ norms for $p>3$ propagated, see [DV00a] and [Wu14] for hard and soft potential cases; For the very soft potential case, the moments and $L^{p}$ norms propagated only known to be local in time or global in time for small initial data, see for example [ALL15].
Concerning the variable spatial kernel $\kappa$, similar results hold on torus $\mathbb{T}^{3}=\mathbb{R}^{3}/{\mathbb{Z}}^{3}$ with $L^{p}_{v}L^{1}_{x}$ replaced by $L^{p}(\mathbb{T}^{3}\times\mathbb{R}^{3})$. The proof will be contained in the upcoming Part II. Notice that $L^{p}(\mathbb{R}^{3};L^{1}(\mathbb{T}^{3}))\subset L^{p}(\mathbb{T}^{3}\times\mathbb{R}^{3})$ since by Minkowski and Hölder inequalities
$\Big{(}\int_{\mathbb{R}^{3}}\Big{(}\int_{\mathbb{T}^{3}}f{\,\rm d}x\Big{)}^{p}{\,\rm d}v\Big{)}^{\frac{1}{p}}\leq\int_{\mathbb{T}^{3}}\Big{(}\int_{\mathbb{R}^{3}}f^{p}{\,\rm d}v\Big{)}^{\frac{1}{p}}{\,\rm d}x\leq\|f\|_{L^{p}(\mathbb{T}^{3}\times\mathbb{R}^{3})}.$
@@ 2.4 Entropy dissipation and Fisher information
In this Subsection, we consider the curves on $\mathbb{R}^{3}\times\mathbb{R}^{3}$ and $\mathbb{T}^{3}\times\mathbb{R}^{3}$ satisfying Assumption 2.2, and the kernel $(A,\kappa)$ satisfying Assumption 2.1. In particular, the spatial kernels $\kappa$ have positive upper and lower bounds
$C^{-1}\leq\kappa\leq C,\quad\text{for some constant }C>0.$
For simplification, we write the domain as $\mathbb{R}^{3}\times\mathbb{R}^{3}$, and unless otherwise stated, it refers to both $\mathbb{R}^{3}\times\mathbb{R}^{3}$ and $\mathbb{T}^{3}\times\mathbb{R}^{3}$.
By using of the cross-product identity (2.5)
$|x|^{2}(y\cdot\Pi_{x^{\perp}}y)=|x\times y|^{2}\quad\forall\,x,\,y\in\mathbb{R}^{3}.$
the entropy dissipation (1.10) can be written as
$\mathcal{D}(f)=\frac{1}{2}\int_{\mathbb{R}^{12}}\kappa ff_{*}|v-v_{*}|^{\gamma}\big{|}(v-v_{*})\times(\nabla_{v}\log f-\nabla_{v_{*}}\log f)\big{|}^{2}{\,\rm d}\eta.$ (2.15)
@@
$|v-v_{*}|^{2}|\tilde{\nabla}\log f|^{2}$ (2.16)
$=|v-v_{*}|^{2}\big{(}\nabla_{v}\log f-\nabla_{v_{*}}\log f_{*}\big{)}\cdot\Pi_{(v-v_{*})^{\perp}}\big{(}\nabla_{v}\log f-\nabla_{v_{*}}\log f_{*}\big{)}$
$=\big{|}(v-v_{*})\times(\nabla_{v}\log f-\nabla_{v_{*}}\log f)\big{|}^{2}.$
@@
$\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f\langle v\rangle^{\gamma}\Big{|}\frac{\nabla_{v}f}{f}\Big{|}^{2}{\,\rm d}x{\,\rm d}v\quad\text{and}\quad\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f\langle v\rangle^{\gamma}\Big{|}\frac{\nabla_{v}f}{f}\times v\Big{|}^{2}{\,\rm d}x{\,\rm d}v.$
@@
Let $f\in\mathcal{P}\cap L\log L\cap\mathcal{S}(\mathbb{R}^{3}\times\mathbb{R}^{3})$. Let $\mathcal{D}(f)$ be the entropy dissipation defined as in (2.15), where $\gamma\in(-3,1]$ and the kernels satisfy
$A(v-v_{*})=|v-v_{*}|^{2+\gamma}\quad\text{and}\quad\kappa\geq C_{\kappa}^{-1}$
for some constant $C_{\kappa}>0$. The following bounds hold
$\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f\langle v\rangle^{\gamma}\Big{|}\frac{\nabla_{v}f}{f}\Big{|}^{2}{\,\rm d}x{\,\rm d}v+\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f\langle v\rangle^{\gamma}\Big{|}\frac{\nabla_{v}f}{f}\times v\Big{|}^{2}{\,\rm d}x{\,\rm d}v\leq C(f)\big{(}1+\mathcal{D}(f)\big{)}.$ (2.17)
  • Let $\gamma\in(-3,0]$. Then we have
    $C(f)\lesssim(a_{0}^{-2}+1)\|f\|_{L^{1}_{0,2+|\gamma|}(\mathbb{R}^{3}\times\mathbb{R}^{3})},$
  • Let $\gamma\in(0,1]$. Let $p>\frac{3}{3-\gamma}$. Then we have
    $C(f)\lesssim\|f\|_{L^{1}_{0,2}(\mathbb{R}^{3}\times\mathbb{R}^{3})}^{6}\big{(}\|\langle v\rangle^{\gamma}f\|_{L^{p}_{v}L^{1}_{x}}+\|f\|_{L^{1}_{0,\gamma}(\mathbb{R}^{3}\times\mathbb{R}^{3})}\big{)}.$
In the above,
$a_{0}(f):=\inf_{w\in S^{2}}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}(v\cdot w)^{2}f(x,v){\,\rm d}x{\,\rm d}v.$
Proof of Lemma 2.6.
The proof follows [Des20, Proposition 4] and [Des16, Corollary 2.7] with appropriate modification according to the spatial kernel $x$.
The quantity $a_{0}(f)>0$ is ensured by $f\in L\log L(\mathbb{R}^{3}\times\mathbb{R}^{3})$.
Without loss of generality, we assume
$\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f{\,\rm d}x{\,\rm d}v=1,\quad\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(x,v)v{\,\rm d}x{\,\rm d}v=0,\quad\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(x,v)|v|^{2}{\,\rm d}x{\,\rm d}v=3.$
Indeed, one can normalise $f$ for example by $\int_{\mathbb{R}^{6}}\big{(}v-\int_{\mathbb{R}^{6}}vf\big{)}f{\,\rm d}x{\,\rm d}v$.
We define $a_{i}=\int_{\mathbb{R}^{6}}f(x,v)v_{i}^{2}{\,\rm d}x{\,\rm d}v$, then $\sum_{i=1}^{3}a_{i}=3$, and $\int_{\mathbb{R}^{6}}f(x,v)v_{i}v_{j}=a_{i}\delta_{ij}$.
We define $\partial_{j}\mathop{=}\limits^{\textrm{def}}\partial_{v_{j}}$ and for $i,j=1,\ldots,3$:
$q_{ij}=q_{ij}^{f}(x,v,y,w)=\Big{(}(v-w)\times\Big{(}\frac{\nabla f}{f}(x,v)-\frac{\nabla f}{f}(y,w)\Big{)}\Big{)}_{i,j}$ (2.18)
$=(v_{i}-w_{i})\Big{(}\frac{\partial_{j}f}{f}(x,v)-\frac{\partial_{j}f}{f}(y,w)\Big{)}-(v_{j}-w_{j})\Big{(}\frac{\partial_{i}f}{f}(x,v)-\frac{\partial_{i}f}{f}(y,w)\Big{)}.$
@@
$\mathcal{D}(f)=\frac{1}{2}\int_{\mathbb{R}^{12}}\kappa(x-y)|v-w|^{\gamma}f(x,v)f(y,w)|q^{f}|^{2}{\,\rm d}\eta.$ (2.19)
  • Let $\gamma\in(-3,0]$. The following proof also works for $\gamma\in(0,1]$. On the other hand, better bounds for hard potential cases will be shown in the second part.
@@
    $\Big{(}v\times\frac{\nabla_{v}f}{f}(x,v)\Big{)}_{ij}-w_{i}\frac{\partial_{i}f}{f}(x,v)+w_{j}\frac{\partial_{j}f}{f}(x,v)$ (2.20)
    $=v_{i}\frac{\partial_{j}f}{f}(y,w)-v_{j}\frac{\partial_{i}f}{f}(y,w)-\Big{(}w\times\frac{\nabla_{v}f}{f}(y,w)\Big{)}_{ij}+q_{ij}.$
    We multiply (2.20) by $f(y,w)$ and integral over $(y,w)\in\mathbb{R}^{3}\times\mathbb{R}^{3}$ to derive
    $\Big{(}v\times\frac{\nabla_{v}f}{f}(x,v)\Big{)}_{ij}-\Big{(}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}w_{i}f(y,w)\Big{)}\frac{\partial_{i}f}{f}(x,v)+\Big{(}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}w_{j}f(y,w)\Big{)}\frac{\partial_{j}f}{f}(x,v)$
    $=\Big{(}v\times\frac{\nabla_{v}f}{f}(x,v)\Big{)}_{ij}$
    $=v_{i}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\partial_{j}f(y,w)-v_{j}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\partial_{i}f(y,w)+\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(y,w)q_{ij}{\,\rm d}y{\,\rm d}w$
    $=\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(y,w)q_{ij}{\,\rm d}y{\,\rm d}w,$
    which implies
    $\Big{(}v\times\frac{\nabla_{v}f}{f}(x,v)\Big{)}_{ij}=\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(y,w)q_{ij}{\,\rm d}y{\,\rm d}w.$ (2.21)
    Next we multiply (2.20) by $w_{j}f(y,w)$ ($i\neq j$) and integral over $(y,w)\in\mathbb{R}^{3}\times\mathbb{R}^{3}$, using the above identity and noticing that
    $\Big{(}v\times\frac{\nabla_{v}f}{f}(x,v)\Big{)}_{ij}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}w_{j}f(y,w){\,\rm d}y{\,\rm d}w=0$
    to derive
    $-\Big{(}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}w_{i}w_{j}f(y,w)\Big{)}\frac{\partial_{i}f}{f}(x,v)+\Big{(}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}w_{j}^{2}f(y,w)\Big{)}\frac{\partial_{j}f}{f}(x,v)$
    $=a_{j}\frac{\partial_{j}f}{f}(x,v)$
    $=v_{i}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}w_{j}\partial_{j}f(y,w)-v_{j}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}w_{j}\partial_{i}f(y,w)+\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}w_{j}f(y,w)q_{ij}{\,\rm d}y{\,\rm d}w$
    $=-v_{i}+\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}w_{j}f(y,w)q_{ij}{\,\rm d}y{\,\rm d}w,$
    which implies
    $a_{j}\frac{\partial_{j}f}{f}(x,v)+v_{i}=\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}w_{j}f(y,w)q_{ij}{\,\rm d}y{\,\rm d}w.$ (2.22)
    Using the relations (2.21) and (2.22), we can bound the Fisher and cross-Fisher information. For the cross-Fisher information, by Cauchy–Schwarz inequality, we have
    $\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(x,v)\Big{|}v\times\frac{\nabla_{v}f}{f}(x,v)\Big{|}^{2}\langle v\rangle^{\gamma}{\,\rm d}x{\,\rm d}v$ (2.23)
    $=\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(x,v)\Big{|}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(y,w)q^{f}{\,\rm d}y{\,\rm d}w\Big{|}^{2}\langle v\rangle^{\gamma}{\,\rm d}x{\,\rm d}v$
    $\leq C_{\kappa}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(x,v)\Big{(}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\kappa(x-y)f(y,w)|q^{f}|^{2}|v-w|^{\gamma}{\,\rm d}y{\,\rm d}w\Big{)}$
    ×$\quad\times\Big{(}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(y,w)|v-w|^{-\gamma}\langle v\rangle^{\gamma}{\,\rm d}y{\,\rm d}w\Big{)}{\,\rm d}x{\,\rm d}v$
    $\leq 2C_{\kappa}\mathcal{D}(f)\sup_{v\in\mathbb{R}^{3}}\langle v\rangle^{\gamma}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(y,w)|v-w|^{-\gamma}{\,\rm d}y{\,\rm d}w$
    $=2C_{\kappa}\mathcal{D}(f)C^{\pm}_{-\gamma,0}(f),$
    where we use the inequality $\kappa\geq C_{\kappa}^{-1}$ (without loss of generality, we assume $C_{\kappa}\geq 1$) and the identity (2.19). The boundedness of the constant $C_{-\gamma,0}(f)>0$ is given in Lemma 2.7 below.
    Similarly, by Cauchy–Schwarz inequality, we have
    $\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(x,v)\Big{|}\frac{\partial_{i}f}{f}(x,v)+\frac{v_{i}}{a_{j}}\Big{|}^{2}\langle v\rangle^{\gamma}{\,\rm d}x{\,\rm d}v$
    $\leq a_{j}^{-2}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(x,v)\Big{|}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(y,w)w_{j}q_{ij}{\,\rm d}y{\,\rm d}w\Big{|}^{2}\langle v\rangle^{\gamma}{\,\rm d}x{\,\rm d}v$
    $\leq C_{\kappa}a_{j}^{-2}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(x,v)\Big{(}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\kappa(x-y)f(y,w)|q_{ij}|^{2}|v-w|^{\gamma}{\,\rm d}y{\,\rm d}w\Big{)}$
    ×$\quad\times\Big{(}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(y,w)|v-w|^{-\gamma}\langle v\rangle^{\gamma}|w|^{2}{\,\rm d}y{\,\rm d}w\Big{)}{\,\rm d}x{\,\rm d}v$
    $\leq 2C_{\kappa}a_{j}^{-2}\mathcal{D}(f)\sup_{v\in\mathbb{R}^{3}}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(y,w)|v-w|^{-\gamma}\langle v\rangle^{\gamma}|w|^{2}{\,\rm d}y{\,\rm d}w$
    $=2C_{\kappa}a_{j}^{-2}\mathcal{D}(f)C^{\pm}_{-\gamma,2}(f).$ (2.24)
    The boundedness of the constant $C^{\pm}_{-\gamma,2}(f)>0$ is given in Lemma 2.7 below.
    Then, by using of the estimate (2.4) and Cauchy-Schwarz inequality, we can bound the Fisher information as follows:
    $\sum_{i=1}^{3}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(x,v)\Big{|}\frac{\partial_{i}f}{f}(x,v)\Big{|}^{2}\langle v\rangle^{\gamma}{\,\rm d}x{\,\rm d}v$
    $\leq\sum_{i,j=1}^{3}2a_{j}^{-2}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(x,v)v_{i}^{2}\langle v\rangle^{\gamma}{\,\rm d}x{\,\rm d}v+2C_{\kappa}a_{j}^{-2}\mathcal{D}(f)C^{\pm}_{-\gamma,2}(f)$
    $\leq 2a_{0}^{-2}(C_{\kappa}+1)\Big{(}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(x,v)|v|^{2}\langle v\rangle^{\gamma}{\,\rm d}x{\,\rm d}v+\mathcal{D}(f)C^{\pm}_{-\gamma,2}(f)\Big{)}.$
    By combining (2.23) and (2.4), the following bounds hold
    $\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f\langle v\rangle^{\gamma}\Big{|}\frac{\nabla_{v}f}{f}\times v\Big{|}^{2}{\,\rm d}x{\,\rm d}v+\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f\langle v\rangle^{\gamma}\Big{|}\frac{\nabla_{v}f}{f}\Big{|}^{2}{\,\rm d}x{\,\rm d}v\leq C(f)\big{(}1+\mathcal{D}(f)\big{)},$
    where
    $C(f)$ $=C^{\pm}_{-\gamma,0}(f)+Ca_{0}^{-2}\big{(}C_{-\gamma,2}(f)+\|f\|_{L^{1}_{0,2+|\gamma|}}\big{)}$
    $\lesssim\|f\|_{L^{1}_{0,|\gamma|}}+Ca_{0}^{-2}\big{(}\|f\|_{L^{1}_{0,2+|\gamma|}}+\|f\|_{L^{1}_{0,2+|\gamma|}}\big{)}.$
    When $\gamma\in(-3,0)$, $C^{+}_{-\gamma,0}$ and $C^{+}_{-\gamma,2}$ are given in (2.27), we arrive at (2.17)
    $C(f)\lesssim\|f\|_{L^{1}_{0,|\gamma|}}+a_{0}^{-2}\|f\|_{L^{1}_{0,2+|\gamma|}}.$
  • Let $\gamma\in(0,1]$.
    We take the cross-Fisher information bounds (2.23)
    $\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(x,v)\Big{|}v\times\frac{\nabla_{v}f}{f}(x,v)\Big{|}^{2}\langle v\rangle^{\gamma}{\,\rm d}x{\,\rm d}v\leq 2C_{\kappa}\mathcal{D}(f)C^{-}_{-\gamma,0}(f),$ (2.25)
    where $C^{-}_{-\gamma,0}$ is given in (2.28) below.
    We closely follow [Des16, Corollary 2.7] to derive the Fisher information bounds. Notice that $q_{ij}$ defined in (2.18) can be written as
    $q_{ij}$ $=\Big{(}v_{i}\frac{\partial_{j}f}{f}(x,v)-v_{j}\frac{\partial_{i}f}{f}(x,v)\Big{)}+w_{j}\frac{\partial_{i}f}{f}(x,v)-w_{i}\frac{\partial_{j}f}{f}(x,v)$ (2.26)
    $\quad-v_{i}\frac{\partial_{j}f}{f}(y,w)+v_{j}\frac{\partial_{i}f}{f}(y,w)+\Big{(}w_{i}\frac{\partial_{j}f}{f}(y,w)-w_{j}\frac{\partial_{i}f}{f}(y,w)\Big{)}.$
@@
    We define the function $\phi(r)=(1+2r)^{-\frac{\gamma}{4}-\frac{1}{2}}$ and we write $\phi=\phi(|w|^{2}/2)=\langle w\rangle^{-\frac{\gamma}{2}-1}$. Notice that $|\phi|\leq 1$, $|\phi^{\prime}(r)|\leq\frac{3}{2}$ and $\nabla\phi=w\phi^{\prime}(|w|^{2}/2)$.

### v2

@@ 2.3 Assumptions on curves
One can make a similar assumption by replacing $L^{p}_{v}L^{1}_{x}$ by $L^{1}_{x}L^{p}_{v}$. The $L^{p}$ bound is to ensure the integrability of $\int_{\mathbb{R}^{4d}}ff_{*}|v-v_{*}|^{\gamma}{\,\rm d}\eta$, for which we use in Lemma 2.10 below, and we make no distinction between $L^{1}_{x}L^{p}_{v}$ and $L^{p}_{v}L^{1}_{x}$ there.
In the follow-up Part II paper, we will show the existence and regularity of the $\mathcal{H}$ solutions to the fuzzy Landau equation (1.1).
Concerning the propagation of moment and $L^{p}_{v}L^{1}_{x}$ norms, we first see the case of the spatial kernel $\kappa=1$. We integrate (1.1) over $x\in\mathbb{R}^{d}$ and define $F(v)=\int_{\mathbb{R}^{d}}f{\,\rm d}x$, which satisfies the homogeneous Landau equations (2.7). The regularity results for homogeneous Landau equations hold: For $\gamma\in[-2,1]$, all moment and $L^{p}$ norms for $p>3$ propagated, see [DV00] and [Wu14] for hard and soft potential cases; For the very soft potential case, the moment and $L^{p}$ norms propagated only known to be local in time or global in time for small initial data, see for example [ALL15].
Concerning the variable spatial kernel $\kappa$, similar results hold on torus $\mathbb{T}^{d}=\mathbb{R}^{d}/{\mathbb{Z}}^{d}$ with $L^{p}_{v}L^{1}_{x}$ replaced by $L^{p}(\mathbb{T}^{d}\times\mathbb{R}^{d})$. The proof will be contained in the upcoming Part II. Notice that $L^{p}(\mathbb{R}^{d};L^{1}(\mathbb{T}^{d}))\subset L^{p}(\mathbb{T}^{d}\times\mathbb{R}^{d})$ since by Minkowski and Hölder inequalities
$\Big(\int_{\mathbb{R}^{d}}\Big(\int_{\mathbb{T}^{d}}f{\,\rm d}x\Big)^{p}{\,\rm d}v\Big)^{\frac{1}{p}}\leq\int_{\mathbb{T}^{d}}\Big(\int_{\mathbb{R}^{d}}f^{p}{\,\rm d}v\Big)^{\frac{1}{p}}{\,\rm d}x\leq\|f\|_{L^{p}(\mathbb{T}^{d}\times\mathbb{R}^{d})}.$
@@ 2.4 Entropy dissipation and Fisher information
In this Subsection, we consider the curves satisfying Assumption 2.7, and the kernel $(A,\kappa)$ satisfying Assumption 2.1.
For simplification, we write the domain as $\mathbb{R}^{2d}$, and unless otherwise stated, it refers to both $\mathbb{R}^{2d}$ and $\mathbb{T}^{d}\times\mathbb{R}^{d}$.
In the subsequent analysis, we will use the following identity
$|x|^{2}(y\cdot\Pi_{x^{\perp}}y)=\frac{1}{2}\sum_{i,j=1}^{d}|x_{i}y_{j}-x_{j}y_{i}|^{2}\quad\forall\,x,\,y\in\mathbb{R}^{d},$
which can be verified by direct computations:
$|x|^{2}(y\cdot\Pi_{x^{\perp}}y)$ $=|x|^{2}\Big[y-\frac{x\otimes x}{|x|^{2}}y\Big]\cdot y$
$=|x|^{2}|y|^{2}-(x\otimes x)y\cdot y$
$=|x|^{2}|y|^{2}-\sum_{i,j}(x\otimes x)_{ij}y_{i}y_{j}$
$=(\sum_{i}x_{i}^{2})(\sum_{j}y_{j}^{2})-\sum_{i,j}x_{i}x_{j}y_{i}y_{j}$
$=(\sum_{i}x_{i}^{2})(\sum_{j}y_{j}^{2})-\Big(\sum_{i}x_{i}y_{i}\Big)^{2}$
$=\frac{1}{2}\sum_{ij}|x_{i}y_{j}-x_{j}y_{i}|^{2}.$
Notice that in dimension three, this reads $|x|^{2}(y\cdot\Pi_{x^{\perp}}y)=|x\times y|^{2}$. To simplify the notation, we write
$[x,y]_{ij}=x_{i}y_{j}-x_{j}y_{i},\quad i,j=1,\dots,d.$
Then the entropy dissipation (1.10) can be written as
$\mathcal{D}(f)=$ $\frac{1}{4}\sum_{i,j=1}^{d}\int_{\mathbb{R}^{4d}}\kappa ff_{*}|v-v_{*}|^{\gamma}|[v-v_{*},\nabla_{v}\log f-\nabla_{v_{*}}\log f_{*}]_{ij}\big|^{2}{\,\rm d}\eta.$ (2.15)
@@
$|\widetilde{\nabla}\log f|^{2}$ (2.16)
$=$ ${}|v-v_{*}|^{\gamma}|v-v_{*}|^{2}\big(\nabla_{v}\log f-\nabla_{v_{*}}\log f_{*}\big)\cdot\Pi_{(v-v_{*})^{\perp}}\big(\nabla_{v}\log f-\nabla_{v_{*}}\log f_{*}\big)$
$=$ ${}\frac{|v-v_{*}|^{\gamma}}{2}\sum_{i,j=1}^{d}\big|[v-v_{*},\nabla_{v}\log f-\nabla_{v_{*}}\log f_{*}]_{ij}\big|^{2}.$
@@
$\int_{\mathbb{R}^{2d}}f\langle v\rangle^{\gamma}\Big|\frac{\nabla_{v}f}{f}\Big|^{2}{\,\rm d}x{\,\rm d}v\quad\text{and}\quad\sum_{i,j=1}^{d}\int_{\mathbb{R}^{2d}}f\langle v\rangle^{\gamma}\Big|\Big[v,\frac{\nabla_{v}f}{f}\Big]_{ij}\Big|^{2}{\,\rm d}x{\,\rm d}v.$
@@
Let $f\in\mathcal{P}\cap L\log L\cap\mathcal{S}(\mathbb{R}^{2d})$. Let $\mathcal{D}(f)$ be the entropy dissipation defined as in (2.15), where $\gamma\in(-\gamma_{d},1]$ and the kernels satisfy
$A(v-v_{*})=|v-v_{*}|^{2+\gamma}.$
The following bounds hold
$\sum_{i,j=1}^{d}\int_{\mathbb{R}^{2d}}f\langle v\rangle^{\gamma}\Big|\Big[\frac{\nabla_{v}f}{f},v\Big]_{ij}\Big|^{2}{\,\rm d}x{\,\rm d}v+\int_{\mathbb{R}^{2d}}f\langle v\rangle^{\gamma}\Big|\frac{\nabla_{v}f}{f}\Big|^{2}{\,\rm d}x{\,\rm d}v$ $\leq C(f)\big(1+\mathcal{D}(f)\big).$ (2.17)
  • Let $\gamma\in(-\gamma_{d},0]$. Then we have
    $C(f)\lesssim\|f\|_{L^{1}_{0,2}(\mathbb{R}^{2d})}^{7}+\|f\|_{L^{1}_{0,|\gamma|}(\mathbb{R}^{2d})}.$
  • Let $\gamma\in(0,1]$. Let $p>\frac{d}{d-\gamma}$. Then we have
    $C(f)\lesssim$ $\|f\|_{L^{1}_{0,2}(\mathbb{R}^{2d})}^{6}\big(\|f\|_{L^{1}_{0,2+\gamma}(\mathbb{R}^{2d})}+\|f\|_{L^{1}_{0,\gamma}(\mathbb{R}^{2d})}$
    $\quad+\min(\|\langle v\rangle^{\gamma}f\|_{L^{p}_{v}L^{1}_{x}},\|\langle v\rangle^{\gamma}f\|_{L^{1}_{x}L^{p}_{v}})\big).$
Proof.
The proof follows [Des16, Corollary 2.7] with appropriate modification according to the spatial kernel $x$.
We define $\partial_{j}\mathop{=}\limits^{\textrm{def}}\partial_{v_{j}}$ and for $i,j=1,\ldots,d$:
$q_{ij}=q_{ij}^{f}(x,v,y,w)=\Big[v-w,\frac{\nabla f}{f}(x,v)-\frac{\nabla f}{f}(y,w)\Big]_{i,j}$ (2.18)
$=(v_{i}-w_{i})\Big(\frac{\partial_{j}f}{f}(x,v)-\frac{\partial_{j}f}{f}(y,w)\Big)-(v_{j}-w_{j})\Big(\frac{\partial_{i}f}{f}(x,v)-\frac{\partial_{i}f}{f}(y,w)\Big)$
$=\Big(v_{i}\frac{\partial_{j}f}{f}(x,v)-v_{j}\frac{\partial_{i}f}{f}(x,v)\Big)+\Big(w_{j}\frac{\partial_{i}f}{f}(x,v)-w_{i}\frac{\partial_{j}f}{f}(x,v)\Big)+\Big(v_{j}\frac{\partial_{i}f}{f}(y,w)-v_{i}\frac{\partial_{j}f}{f}(y,w)\Big)$
$\quad+\Big(w_{i}\frac{\partial_{j}f}{f}(x,v)-w_{j}\frac{\partial_{i}f}{f}(x,v)\Big).$
@@
$\mathcal{D}(f)=\frac{1}{4}\int_{\mathbb{R}^{4d}}\kappa(x-y)|v-w|^{\gamma}f(x,v)f(y,w)|q^{f}|^{2}{\,\rm d}\eta.$ (2.19)
@@
$\Big[v,\frac{\nabla_{v}f}{f}(x,v)\Big]_{ij}-w_{i}\frac{\partial_{i}f}{f}(x,v)+w_{j}\frac{\partial_{j}f}{f}(x,v)$ (2.20)
$=$ ${}v_{i}\frac{\partial_{j}f}{f}(y,w)-v_{j}\frac{\partial_{i}f}{f}(y,w)-\Big(w_{i}\frac{\partial_{j}f}{f}(y,w)-w_{j}\frac{\partial_{i}f}{f}(y,w)\Big)+q_{ij}.$
@@
To obtain a better bound, we define the weight function $\phi(r)=(1+2r)^{-\frac{|\gamma|}{4}-\frac{1}{2}}$ and we write $\phi=\phi(|w|^{2}/2)=\langle w\rangle^{-\frac{|\gamma|}{2}-1}$. Notice that $|\phi|\leq 1$, $|\phi^{\prime}(r)|\leq\frac{3}{2}$ and $\nabla\phi=w\phi^{\prime}(|w|^{2}/2)$.